\begin{thebibliography}{12}
\bibitem{Ak}  N.~I. Akhiezer,  {\it The Classical Moment Problem and Some Related
 Questions in Analysis}. English translation, Oliver and Boyd, Edinburgh,
 1965.


\bibitem{B:C} C.~Berg and J.~P.~R.~Christensen,   Density questions in the classical theory of moments,  Ann. Inst. Fourier  {\bf 31}, no. 3 (1981), 99--114.

\bibitem{B:S1} C.~Berg and R.~Szwarc,  The Smallest Eigenvalue of Hankel Matrices, Constr. Approx. {\bf 34} (2011), 107--133.

\bibitem{B:S2} C.~Berg and R.~Szwarc,  Inverse of infinite Hankel moment matrices, SIGMA {\bf 14} (2018), 109, 48 pages.

\bibitem{B:S3} C.~Berg and R.~Szwarc,  Closable Hankel Operators and Moment Problems,
Integr. Equ. Oper. Theory {\bf 92}(1) (2020), 1--9.

\bibitem{B:S4} C.~Berg and R.~Szwarc, Self-adjoint operators associated with Hankel moment matrices, Journal of Functional Analysis, {\bf 283} (2022), 109674. 

\bibitem{B:S5} C.~Berg and R.~Szwarc, Indeterminate Jacobi operators, J. Operator Theory
93:1 (2025), 101--123.


\bibitem{Pe} H.~L.~Pedersen, The Nevanlinna matrix of entire functions associated with a shifted indeterminate Hamburger moment problem, Math. Scand. {\bf 74} (1994), 152--160.

\bibitem{Ri} M.~Riesz, Sur le probl{\`e}me des moments et le th{\'e}or{\`e}me de Parseval correspondant, Acta Litt. Ac. Sci. Szeged {\bf 1} (1923), 209--225.

\bibitem{S} B.~Simon, The classical moment problem as a self-adjoint finite difference operator, Adv. Math. {\bf 137} (1998), 82--203.

\bibitem{Sch} K.~Schm\"{u}dgen, {\it The Moment Problem}, Graduate Texts in Mathematics Vol. 277. Springer International Publishing AG 2017. 

\end{thebibliography}

\documentclass [12pt]{article}
\usepackage{amssymb,amsmath,amsthm}
\usepackage{color}

\DeclareMathOperator{\spn}{span}

\usepackage[cp1250]{inputenc}
\usepackage[active]{srcltx}

\newtheorem{thm}{Theorem}[section]
\newtheorem{prop}[thm]{Proposition}
\newtheorem{lem}[thm]{Lemma}
 \newtheorem{rem}[thm]{Remark}
 
\newtheorem{defn}[thm]{Definition}

\DeclareMathOperator{\supp}{supp}
\DeclareMathOperator{\ext}{ext}
\newcommand{\C}{\mathbb{C}}
\newcommand{\R}{\mathbb{R}}

\newcommand{\la}{\lambda}

\long\def\comment#1{{}}
\def\ph{\phantom{-}}
\def\p{\ph}
 \title{Indeterminate Jacobi operators II}
 \author{Christian Berg and Ryszard Szwarc}

\begin{document}

 \maketitle
file: besz10.tex

\begin{abstract}  We consider the Jacobi operator $(T,D(T))$ associated with an indeterminate Hamburger moment problem, and present countable subsets $S$ of the domain $D(T)$  such that $\spn(S)$ is dense in $\ell^2$. As an example  we have $S=\{(p_n(u))+B(u)(p_n(0)) \mid D(u)=0\}$, where $(p_n)$ denotes the orthonormal polynomials of the moment problem and $B, D$ are two of the Nevanlinna functions. 
It is also proved that sets like $S$ are optimal in the sense that if one vector is removed, then the span is no longer dense.
\end{abstract}

{\bf Mathematics Subject Classification}: Primary 47B25, 47B36, 44A60

{\bf Keywords}. Jacobi matrices and operators, indeterminate moment problems.

\section{Introduction and main results} This paper is a continuation of \cite{B:S5}  by improving certain results about the domain of the indeterminate Jacobi operator. We  consider the Jacobi matrix $J$ associated with a moment sequence $s=(s_n)_{n\ge 0}$ of the form
\begin{equation}\label{eq:mom}
s_n=\int x^n\,d\mu(x),\quad n=0,1,\ldots,
\end{equation}
where $\mu$ is a positive measure on $\R$  with infinite support and moments of every order. It is a tridiagonal matrix of the  form
\begin{equation}\label{eq:Jac}
J=\begin{pmatrix}
b_0 & a_0 & 0 & \hdots\\
a_0 & b_1 & a_1 & \hdots\\
0 & a_1 & b_2 & \hdots\\
\vdots &\vdots & \vdots & \ddots
\end{pmatrix},
\end{equation}
where $a_n>0, b_n\in\R, n\ge 0$  are given by the three term recurrence relation
\begin{equation*}\label{eq:3term}
xp_n(x)=a_np_{n+1}(x)+b_np_n(x)+a_{n-1}p_{n-1}(x), n\ge 0, \quad a_{-1}:=0.
\end{equation*}
Here $(p_n)_{n\ge 0}$ is the sequence of orthonormal polynomials associated with $\mu$, hence satisfying
\begin{equation*}\label{eq:op}
\int p_n(x)p_m(x)\,d\mu(x)=\delta_{n,m}.
\end{equation*}
As usual $p_n$ is a real polynomial of degree $n$ with positive leading coefficient. As in \cite{B:S5} we follow the terminology of \cite{Sch}. Basic results about the classical moment problem can also be found in \cite{Ak} and \cite{S}. Recent  results about indeterminate moment problems can be found in \cite{B:S1}, \cite{B:S2} \cite{B:S3}, \cite{B:S4}, \cite{B:S5}.

 It is easy to see that the proportional measures $\la\mu, \la>0$ lead to the same Jacobi matrix $J$, and the well-known Theorem of Favard (see \cite[Theorem 5.14]{Sch}) states that any matrix of the form \eqref{eq:Jac} with $a_n>0, b_n\in\R$ comes from a unique  moment sequence  $(s_n)$ as above, normalized such that $s_0=1$. In the following we shall always assume that this normalization holds, and consequently the solutions $\mu$ of \eqref{eq:mom} are probability measures  and $p_0=1$.

The Jacobi matrix acts as a symmetric operator  in the Hilbert space $\ell^2$ of square summable complex sequences. Its domain $\mathcal F$ consists of the complex sequences $(c_n)_{n\ge 0}$ with only finitely many non-zero terms, and the action is multiplication of the matrix $J$ by $c\in\mathcal F$ considered as a column, i.e.,
\begin{equation}\label{eq:J} 
(Jc)_n:=a_{n-1}c_{n-1}+b_n c_n+a_n c_{n+1},\quad n\ge 0.
\end{equation}
Denoting $(e_n)_{n\ge 0}$ the standard orthonormal basis of $\ell^2$, we have
\begin{equation*}\label{eq:F}
\mathcal F=\spn \{e_n|n\ge 0\}.
\end{equation*}

\begin{defn} The Jacobi operator associated with $J$ is by definition the closure $(T,D(T))$ of the symmetric operator $(J,\mathcal F)$.
\end{defn}

It is a classical fact that the closed symmetric operator $(T,D(T))$ has deficiency indices  either $(0,0)$  or $(1,1)$. These cases occur precisely if the  moment sequence \eqref{eq:mom} is {\it determinate} or {\it indeterminate}, i.e., there is exactly one or several solutions $\mu$ satisfying \eqref{eq:mom}.   

 By definition $D(T)$ consists of those $c\in\ell^2$ for which there exists a sequence
 $(c^{(k)})\in\mathcal F$ such that $\lim_{k\to\infty}c^{(k)}=c$ and $(Jc^{(k)})$ is a convergent sequence in $\ell^2$. For  such $c$ we have $Tc=\lim_{k\to\infty} Jc^{(k)}$, and this limit is independent of the choice of approximating sequence $(c^{(k)})$. 
 
 Clearly, $D(T)$ is closed under complex conjugation and 
 $$
 T\overline{c}=\overline{Tc},\quad c\in D(T).
 $$
 
 We recall that the adjoint operator $(T^*,D(T^*))$ is the maximal operator associated with $J$, cf. \cite[Proposition 6.5]{Sch}. In fact, the matrix product of $J$  and any column vector $c$ makes sense, cf. \eqref{eq:J}, and $D(T^*)$ consists of those $c\in\ell^2$ for which the product $Jc$ belongs to $\ell^2$. For $c\in D(T^*)$ we have $T^*c=Jc$.
 
In this paper we will only consider the  indeterminate case of the Jacobi operator $(T,D(T))$, where it is known that  the set of solutions $\mu$ to \eqref{eq:mom} is an infinite convex set $V$. The polynomials of the second kind $(q_n)$ are given as
\begin{equation*}
q_n(z)=\int \frac{p_n(z)-p_n(x)}{z-x}\,d\mu(x), \quad z\in \C,
\end{equation*}
where $\mu\in V$ is arbitrary.

We define  the sequences
\begin{equation}\label{eq:frak}
 \mathfrak{p}_z:=(p_n(z)), \mathfrak{q}_z:=(q_n(z)),\quad z\in\C,
 \end{equation}
 where we have followed the terminology of \cite{Sch}. It is known that they belong to
$\ell^2$ because of indeterminacy, and
   $||\mathfrak{p}_z||$ and $||\mathfrak{q}_z||$ are positive continuous functions for $z\in\C$.  It is therefore possible 
for $c\in\ell^2$ to define entire functions $F_c, G_c$ as
\begin{equation}\label{eq:FGc}
F_c(z)=\sum_{n=0}^\infty c_np_n(z),\quad G_c(z)=\sum_{n=0}^\infty c_nq_n(z),\quad z\in\C.
\end{equation}

It is well known that $F_c\in L^2(\mu)$ for any solution $\mu\in V$ and that
$$
\lim_{n\to\infty} \sum_{k=0}^n c_kp_k(z)= F_c(z)
$$
locally uniformly in $z\in\C$ and in $L^2(\mu)$ for any $\mu\in V$. Furthermore, Parseval's equation holds
\begin{equation}\label{eq:Pars}
\int |F_c(x)|^2\,d\mu(x)=||c||^2,\quad c\in\ell^2, \mu\in V.
\end{equation}

We recall the following four entire functions of two complex variables, called the {\it Nevanlinna functions} of the indeterminate moment problem:
\begin{eqnarray}
A(u,v)&=&(u-v)\sum_{k=0}^\infty q_k(u)q_k(v)\label{eq:A}\\
B(u,v)&=&-1+(u-v)\sum_{k=0}^\infty p_k(u)q_k(v) \label{eq:B}\\
C(u,v)&=&1+(u-v)\sum_{k=0}^\infty q_k(u)p_k(v)\label{eq:C}\\
D(u,v)&=&(u-v)\sum_{k=0}^\infty p_k(u)p_k(v)\label{eq:D},
\end{eqnarray}
see Section 7.1 in \cite{Sch}. They satisfy the  fundamental determinant equation
\begin{equation}\label{eq:2vardet}
A(u,v)D(u,v)-B(u,v)C(u,v)=1, \quad u,v\in\C.
\end{equation}
 
We define entire functions of one variable by setting the second variable to 0, i.e., 
\begin{equation}\label{eq:A-D}
A(u)=A(u,0),\;B(u)=B(u,0),\;C(u)=C(u,0),\;D(u)=D(u,0),
\end{equation}
and \eqref{eq:2vardet} becomes

\begin{equation}\label{eq:det}
A(u)D(u)-B(u)C(u)=1,\quad u\in\C.
\end{equation}

By Section 6.5 in \cite{Sch} we have
\begin{equation}\label{eq:p,q}
\mathfrak{p}_z, \mathfrak{q}_z \in D(T^*),\quad T^*\mathfrak{p}_z=z\mathfrak{p}_z, 
T^*\mathfrak{q}_z=e_0+z\mathfrak{q}_z,\quad z\in\C.
\end{equation}

A main result of \cite{B:S5} states the following:

\begin{thm}\label{thm:pqmain} For all $z\in\C$ we have $\mathfrak{p}_z, \mathfrak{q}_z\notin D(T)$.

Let $u,v\in\C$ be given.
\begin{enumerate}
\item[(i)] There exists $\alpha\in\C$ such that $\mathfrak{p}_u+\alpha \mathfrak{p}_v\in D(T)$ if and only if $D(u,v)=0$. In the affirmative case $\alpha$  is uniquely determined as
$\alpha=B(u,v)$. 

\item[(ii)] There exists $\beta\in\C$ such that $\mathfrak{q}_u+\beta \mathfrak{q}_v\in D(T)$ if and only if $A(u,v)=0$. In the affirmative case $\beta$  is uniquely determined as $\beta=-C(u,v)$.

\item[(iii)] There exists $\gamma\in\C$ such that $\mathfrak{p}_u+\gamma \mathfrak{q}_v\in D(T)$ if and only if $B(u,v)=0$. In the affirmative case $\gamma$  is uniquely determined as $\gamma=-D(u,v)$. In particular $\mathfrak{p}_u+\gamma\mathfrak{q}_u\notin D(T)$ for all $u, \gamma\in\C$.
\end{enumerate}
\end{thm}

From  this theorem we have the following concrete subspaces of $D(T)$:
\begin{eqnarray*}
P&=&\spn\{\mathfrak{p}_u+B(u,v){\mathfrak{p}}_v \mid u,v\in\C, D(u,v)=0, u\neq v\},\\
Q&=&\spn\{\mathfrak{q}_u-C(u,v){\mathfrak{q}}_v \mid u,v\in\C, A(u,v)=0, u\neq v\},\\
M&=&\spn\{\mathfrak{p}_u-D(u,v){\mathfrak{q}}_v \mid u,v\in\C, B(u,v)=0\}.
\end{eqnarray*}
Note that $\mathfrak{p}_u+B(u,v)\mathfrak{p}_{v} =\mathfrak{q}_u-C(u,v)\mathfrak{q}_v=0$  for $u=v$.

For a fixed number $v_0\in\R$ we define the following subspaces of $P,Q,M$
respectively
\begin{eqnarray}
P(v_0)&=&\spn\{\mathfrak{p}_u+B(u,v_0)\mathfrak{p}_{v_0} \mid u\in\R, D(u,v_0)=0, u\neq v_0\},\label{eq:P}\\
Q(v_0)&=&\spn\{\mathfrak{q}_u-C(u,v_0){\mathfrak{q}}_{v_0} \mid u\in\R, A(u,v_0)=0, u\neq v_0\},\label{eq:Q}\\
M(v_0)&=&\spn\{\mathfrak{p}_u-D(u,v_0){\mathfrak{q}}_{v_0} \mid u\in\R, B(u,v_0)=0\}\label{eq:M}.
\end{eqnarray}
In these definitions it is important to remember, that if $F$ is any of the functions $A,B,C,D$ of two variables, then
$$
Z(F)_{v_0}:=\{u\in\C \mid F(u,v_0)=0\}
$$
is a countably infinite set of real numbers, cf. Theorem 1.3 i \cite{B:S5}.

Letting $\mu[v_0]$ denote the unique N-extremal measure in $V$ with $v_0\in\supp(\mu[v_0])$,
cf. Proposition~\ref{thm:supportNext}, we have by Theorem 3 in \cite{B:C}
\begin{equation}\label{eq:supp}
\supp(\mu[v_0])=\{u\in\R \mid D(u,v_0)=0\},
\end{equation}
and hence
\begin{equation} 
P(v_0)=\spn\{\mathfrak{p}_u+B(u,v_0)\mathfrak{p}_{v_0} \mid u\in\supp(\mu[v_0])
\setminus\{v_0\}\}.
\end{equation}

In the next section we recall the parametrization $\mu_t, t\in\R^*$ of the N-extremal measures in $V$, and by Proposition~\ref{thm:supportNext} we have
\begin{equation}\label{eq:N-ext}
\mu[v_0]=\mu_t,\quad t=-B(v_0)/D(v_0)
\end{equation}
with the convention that $t=\infty$ if $D(v_0)=0$.

 The following result sharpens that $D(T)$ is dense in $\ell^2$ by giving concrete dense subspaces of $D(T)$, which are all spanned by countably many vectors. It is the first main result of this paper.

  \begin{thm}\label{thm:concrete} For each $v_0\in\R$ the subspaces $P(v_0), Q(v_0), M(v_0)$ are dense in $\ell^2$.
\end{thm}

\begin{rem}\label{thm:optimal}
{\rm Let $u_1\neq u_2$ satisfy $D(u_1,v_0)=D(u_2,v_0)=0$, where $v_0\in\R$. Then $\mathfrak{p}_{u_1},\mathfrak{p}_{u_2}$ are orthogonal in $\ell^2$ because
$$
\langle \mathfrak{p}_{u_1},\mathfrak{p}_{u_2}\rangle=\frac{D(u_1,u_2)}{u_1-u_2},
$$
and by Theorem 5.1 in \cite{B:S5}
$$
D(u_1,u_2)=D(u_1,v_0)C(v_0,u_2)-B(u_1,v_0)D(v_0,u_2)=0.
$$
In particular, if $u_1\neq v_0$ satisfies $D(u_1,v_0)=0$, then
$$
\langle \mathfrak{p}_{u_1}, \mathfrak{p}_u+B(u,v_0)\mathfrak{p}_{v_0}\rangle=0 \;\mbox{for}\; D(u,v_0)=0, u\neq u_1,
$$
showing that $\mathfrak{p}_{u_1}$ is orthogonal to
$$
\spn\{\mathfrak{p}_u+B(u,v_0)\mathfrak{p}_{v_0} \mid D(u,v_0)=0,u\neq u_1\},
$$
so the latter cannot be dense in $\ell^2$.
In other  words, the family \eqref{eq:P} is optimal for density in $\ell^2$.

Similarly the family \eqref{eq:Q} is optimal for density in $\ell^2$.
}
\end{rem}

The proof of Theorem~\ref{thm:concrete} will be given in Section 3.

\section{Preliminaries about indeterminate moment problems}

For the proof of Theorem~\ref{thm:concrete} we need the following polynomial approximations to the Nevanlinna functions.
 
\begin{prop}\cite[Proposition 5.24]{Sch}\label{thm:An-Dn} For $u,v\in\C$ and $n\ge 0$ we have
\begin{eqnarray*}
A_n(u,v)&:=&(u-v)\sum_{k=0}^nq_k(u)q_k(v)=
a_n\left|\begin{array}{cc}
 q_{n+1}(u)&\;q_{n+1}(v)\\q_{n}(u)&\;q_{n}(v)\end{array}\right|\\
B_n(u,v)&:=&-1+(u-v)\sum_{k=0}^n p_k(u)q_k(v)=
a_n\left|\begin{array}{cc}
p_{n+1}(u)&\;q_{n+1}(v)\\p_{n}(u)&\;q_{n}(v)\end{array}\right|\\
C_n(u,v)&:=&1+(u-v)\sum_{k=0}^n q_k(u)p_k(v)=
a_n\left|\begin{array}{cc}
q_{n+1}(u)&\;p_{n+1}(v)\\
q_{n}(u)&\;p_{n}(v)\end{array}\right|\\
D_n(u,v)&:=&(u-v)\sum_{k=0}^np_k(u)p_k(v)=
a_n\left|\begin{array}{cc}
p_{n+1(}u)&\;p_{n+1}(v)\\
p_{n}(u)&\;p_{n}(v)\end{array}\right|.
\end{eqnarray*}
\end{prop}

The Jacobi operator $(T,D(T))$ has deficiency indices $(1,1)$ and the self-adjoint extensions in
$\ell^2$ can be parametrized as the operators $T_t, t\in\R^*=\R\cup \{\infty\}$ with domain  
\begin{equation}\label{eq:domTt}
D(T_t)=D(T)\oplus \C  (\mathfrak{q}_0 + t\mathfrak{p}_0) \;\mbox{for}\;t\in\R,\quad
D(T_\infty)=D(T)\oplus \C \mathfrak{p}_0
\end{equation}
and defined by the restriction of $T^*$ to the domain, cf. \cite[Theorem 6.23]{Sch}.
We recall that $\mathfrak{p}_0, \mathfrak{q}_0$ are defined in \eqref{eq:frak}.

For $t\in\R^*$ we define the solutions to the moment sequence \eqref{eq:mom}
\begin{equation}\label{eq:Next}
\mu_t(\cdot):=\langle E_t(\cdot)e_0,e_0\rangle,
\end{equation}
where $E_t(\cdot)$ is the spectral measure of the self-adjoint operator $T_t$.

The measures $\mu_t, t\in\R^*$ are precisely those measures $\mu\in V$  for which the polynomials $\C[x]$ are dense in $L^2(\mu)$ according to a famous theorem of M. Riesz, cf. \cite{Ri}.
 They are called N-extremal  in \cite{Ak} and von Neumann solutions in \cite{S}, and they form a compact subset of the set $\ext(V)$ of extreme points of the convex set $V$. However, $\ext(V)$ is known to be a dense subset of $V$.  The N-extremal measures are characterized by the formula
\begin{equation}\label{eq:Npar}
\int\frac{d\mu_t(x)}{x-z}=-\frac{A(z)+tC(z)}{B(z)+tD(z)},\quad z\in \C\setminus\R, t\in\R^*,
\end{equation}
where $A,\ldots,D$ are the entire functions given in \eqref{eq:A-D}, cf. \cite[Theorem 7.6]{Sch}. Recall that \eqref{eq:det} holds, so the right-hand side of \eqref{eq:Npar} is a M\"{o}bius transformation in $t$. We note in passing that the solutions $\mu\in V$ different from the N-extremal ones are given in \eqref{eq:Npar}, when $t$ is replaced by a non-degenerate Pick function $\varphi:\C\setminus\R \to\C$, cf. Theorem 7.13 in \cite{Sch}.

We summarize some of the properties of $\mu_t$, which can be found in \cite{Ak} and \cite{Sch}.

\begin{prop}\label{thm:supportNext}
\begin{enumerate}
\item[(i)] The  solution $\mu_t$ is a discrete measure with support equal to the countable zero set $\Lambda_t$ of the entire function $B(z)+tD(z)$, with the convention that $\Lambda_\infty$ is the zero set of $D$. We have $\Lambda_t\subset\R$ for $t\in\R^*$.
\item[(ii)] The support of two different N-extremal solutions are disjoint and interlacing. Each point $x_0\in\R$ belongs to the support of a unique N-extremal measure $\mu_t$, where
$t\in\R^*$ is given  as $t=-B(x_0)/D(x_0)$ if $D(x_0)\neq 0$ and $t=\infty$ if $D(x_0)=0$.
\item[(iii)] If $x_0\in\R$ belongs to the support of the N-extremal measure $\mu_t$, then the measure $\mu_t-\mu_t(x_0)\delta_{x_0}$ is determinate.
\end{enumerate}
\end{prop}

Putting $z=0$ in \eqref{eq:Npar}, which is possible when $0\notin \supp(\mu_t)$, leads to
\begin{equation}\label{eq:Npar0}
\int\frac{d\mu_t(x)}{x}=t,\quad t\in\R.
\end{equation}

\section{Proof of Theorem~\ref{thm:concrete}}
 In the proof of Theorem~\ref{thm:concrete} we need the following result about the functions $B(u,v), D(u,v)$. It is  of independent interest.

\begin{lem}\label{thm:1-2} (i) For $\mu\in V, v\in\C$ and any polynomial $p$ the functions $p(u)B(u,v), p(u)D(u,v)$ belong to $L^1(\mu)$ as functions of $u$ and
$$
\int p(u)B(u,v)\,d\mu(u)=\int p(u)D(u,v)\,d\mu(u)=0.
$$ 

(ii) For $v_0\in\R$ we have $B(u,v_0)\notin L^2(\mu[v_0])$.

(iii) For $v_0\in\R$ we have $D(u,v_0)\notin L^2(\mu)$ for any N-extremal measure $\mu\neq \mu[v_0]$, while $D(u,v_0)$ is the zero element in $L^2(\mu[v_0])$.
 \end{lem} 

\begin{proof} (i). With the notation of \eqref{eq:FGc} we have
\begin{eqnarray*}
p(u)B(u,v)&=&-p(u)+p(u)(u-v)F_{(q_k(v))}(u),\\
 p(u)D(u,v)&=&p(u)(u-v)F_{(p_k(v))}(u),
\end{eqnarray*}
so it is clear from the Cauchy-Schwarz inequality that $p(u)B(u,v), p(u)D(u,v)\in L^1(\mu)$ as functions of $u$ and furthermore that
$p(u)B_n(u,v)\to p(u)B(u,v)$ and $p(u)D_n(u,v)\to p(u)D(u,v)$  in $L^1(\mu)$ for $n\to\infty$. However,
\begin{equation}\label{eq:int0}
\int p(u)B_n(u,v)\,d\mu(u)=\int p(u)D_n(u,v)\,d\mu(u)=0
\end{equation}
 for $n>\deg(p)$ by Proposition~\ref{thm:An-Dn}, and (i) follows.

(ii). If $B(u,v_0)\in L^2(\mu[v_0])$ it follows from \eqref{eq:int0} that $B(u,v_0)=0$ for all $u\in\supp(\mu[v_0])$ and in particular for $u=v_0$, which is a contradiction since $B(v_0,v_0)=-1$.

(iii). If $D(u,v_0)\in L^2(\mu)$ for an N-extremal measure $\mu$, we get that $D(u,v_0)=0$ for all $u\in\supp(\mu)$. This is not possible if $\mu\neq \mu[v_0]$ because then  
$\supp(\mu)$ is disjoint from $\supp(\mu[v_0])$ and this contradicts \eqref{eq:supp}, which also shows that  $D(u,v_0)$ is the zero element in $L^2(\mu[v_0])$.
\end{proof}

\medskip
{\it Proof of Theorem~\ref{thm:concrete}} (i). Assume that $c\in\ell^2$ is orthogonal to $P(v_0)$ and let us prove that $c=0$.  The orthogonality can be expressed as
$$
F_c(u)+B(u,v_0)F_c(v_0)=0
$$
for all $u\in\supp(\mu[v_0])\setminus\{v_0\}$,
but this equation  clearly holds for $u=v_0$ as well. 

If $F_c(v_0)\neq 0$ then $B(u,v_0)=-F_c(u)/F_c(v_0)$ for $u\in\supp(\mu[v_0])$ and in particular
$B(u,v_0)\in L^2(\mu[v_0])$, which contradicts Lemma~\ref{thm:1-2}. Therefore $F_c(v_0)=0$ and then $F_c(u)=0$ for all $u\in\supp(\mu[v_0])$, hence
$$
0=\int|F_c(u)|^2\,d\mu[v_0](u)=||c||^2
$$
and therefore $c=0$.

(ii). This case  can be deduced from case (i) by using the observation that
the polynomials $(q_{n+1}(x)/q_1(x))_{n\ge 0}$ are the orthonormal polynomials associated with the truncated Jacobi matrix $J^{(1)}$ obtained from $J$ by removing the first row and column. See \cite[p. 28]{Ak}, \cite[p. 122]{B:S5} and \cite{Pe} for details.

(iii). Assume that $c\in\ell^2$ is orthogonal to $M(v_0)$ and let us prove that $c=0$.  The orthogonality can be expressed as
\begin{equation}\label{eq:ortM}
F_c(u)-D(u,v_0)G_c(v_0)=0\;\mbox{for all}\; u\in\R\;\mbox{such that}\; B(u,v_0)=0.
\end{equation} 

From formula (5.6) in \cite{B:S5} we have $B(u,v_0)=0$ if and only if 
\begin{equation*}
 \left\{\begin{array}{cl}
 B(u)-\frac{A(v_0)}{C(v_0)}D(u)=0 & \mbox{if} \;\; C(v_0)\neq 0,\\
 D(u)=0 & \mbox{if}\;\; C(v_0)=0.
 \end{array}
  \right.
  \end{equation*}
From Proposition~\ref{thm:supportNext} this set of $u$'s is the support of the N-extremal measure
$\mu_{t_0}$, where $t_0=-A(v_0)/C(v_0)$, interpreted as $t_0=\infty$ if $C(v_0)=0$.

By formula (5.8) in \cite{B:S5} the orthogonality condition  \eqref{eq:ortM} can  be expressed
\begin{equation}\label{eq:orth}
F_c(u)-(B(u)D(v_0)-D(u)B(v_0))G_c(v_0)=0,\quad u\in\supp(\mu_{t_0}).
\end{equation} 

{\bf Case  $C(v_0)=0$:}

Then $t_0=\infty$ and $\supp(\mu_\infty)=\{u\mid D(u)=0\}$, so \eqref{eq:orth} states
$$
F_c(u)-B(u)D(v_0)G_c(v_0)=0,\quad u\in\supp(\mu_{\infty}).
$$
If $G_c(v_0)\neq 0$ then
$$
B(u)=\frac{F_c(u)}{D(v_0)G_c(v_0)},\quad u\in\supp(\mu_\infty),
$$
but since $0\in\supp(\mu_\infty)$ this contradicts (ii) of Lemma~\ref{thm:1-2}.
Therefore $G_c(v_0)=0$ and from \eqref{eq:Pars} with $\mu=\mu_\infty$, we get that $c=0$.

\medskip
{\bf Case $C(v_0)\neq 0$:}

Then $t_0=-A(v_0)/C(v_0)\in\R$ and we have for $u\in\supp(\mu_{t_0})$ that
$B(u)=-t_0D(u)$  and hence
$$
B(u)D(v_0)-D(u)B(v_0)=D(u)(-t_0D(v_0)-B(v_0))=\frac{D(u)}{C(v_0)},
$$
where we used \eqref{eq:det}.
Equation \eqref{eq:orth} can now be stated
$$
F_c(u)-\frac{D(u)}{C(v_0)}G_c(v_0),\quad u\in\supp(\mu_{t_0}).
$$
If $G_c(v_0)\neq 0$ then $D(u)\in L^2(\mu_{t_0})$, which contradicts
(iii) of Lemma~\ref{thm:1-2} because $D(u)=D(u,0)$ and $0\notin \supp(\mu_{t_0})$.
Therefore $G_c(v_0)=0$ so $F_c(u)=0$ on $\supp(\mu_{t_0})$ and finally $c=0$.
$\square$
\medskip
                  
In Remark~\ref{thm:optimal} we noticed that for fixed $v_0\in\R$, the  family of vectors
$$
\{\mathfrak{p}_u \mid u\in\R, D(u,v_0)=0\}
$$
are mutually orthogonal in $\ell^2$. We claim that the corresponding normalized vectors
\begin{equation}\label{eq:nb}
\tilde{\mathfrak{p}}_u=\mathfrak{p}_u/||\mathfrak{p}_u||
\end{equation}
 form an orthonormal basis in $\ell^2$. In fact, if $c\in\ell^2$ is orthogonal to these vectors, we know that the entire function $F_c$ defined in \eqref{eq:FGc} satisfies $F_c(u)=0$ for $u\in\supp(\mu[v_0])$, which by 
\eqref{eq:Pars} implies that $c=0$.

We next recall the following easily established Lemma:

\begin{lem}\label{thm:dense} Let $x_n, n\ge 1$ be an orthonormal basis of a complex Hilbert space $\mathcal H$ and let $(a_n)_{n\ge 2}$ be a sequence of complex numbers. Then the subspace $\spn \{x_n+a_nx_1, n\ge 2\}$ is dense in $\mathcal H$ if and only if
$\sum |a_n|^2=\infty$.
\end{lem}

Applying  the lemma to the orthonormal basis \eqref{eq:nb}, we get that
the span of the family
$$
||\mathfrak{p}_u||^{-1}(\mathfrak{p}_u+B(u,v_0)\mathfrak{p}_{v_0})=\tilde{\mathfrak{p}}_u+B(u,v_0)\frac{||\mathfrak{p}_{v_0}||}{||\mathfrak{p}_u||}\tilde{\mathfrak{p}}_{v_0}
,\quad D(u,v_0)=0,u\neq v_0,
$$
is dense in $\ell^2$ if and only if
$$
\sum_{u\in\supp(\mu[v_0])\setminus\{v_0\}}\frac{B(u,v_0)^2}{||\mathfrak{p}_u||^2}=\infty.
$$
However, since $\mu[v_0](\{u\})=1/||\mathfrak{p}_u||^2$ for $u\in\supp(\mu[v_0])$, this is equivalent to
$$
\int B(u,v_0)^2\,d\mu[v_0](u)=\infty.
$$
This gives another proof of the first part of Theorem~\ref{thm:concrete} stating that $P(v_0)$ is dense in $\ell^2$ based on Lemma~\ref{thm:1-2} (ii).

\noindent
Christian Berg\\
Department of Mathematical Sciences, University of Copenhagen\\
Universitetsparken 5, DK-2100 Copenhagen, Denmark\\
e-mail: {\tt{berg@math.ku.dk}}

\vspace{0.4cm}
\noindent
Ryszard Szwarc\\
Institute of Mathematics, University of Wroc{\l}aw\\
pl.\ Grunwaldzki 2/4, 50-384 Wroc{\l}aw, Poland\\ 
e-mail: {\tt{szwarc2@gmail.com}}

\end{document}